# OpenVO: Open-World Visual Odometry with Temporal Dynamics Awareness

### v1

@@ 2 Related Work
Geometric Prior Models. Geometric priors such as metric depth and camera intrinsics are crucial for 3D reconstruction and improving scale consistency in visual odometry [21, 55, 41]. Monocular foundation metric depth models [56, 16, 4, 37, 36, 36, 53, 54, 3] enable absolute-scale depth prediction with strong zero-shot generalization across divese domains through large-scale training. Among these, some rely on known camera parameters [56, 16], while others remove this assumption and even achieve competitive performance in estimating focal lengths from monocular observation [4]. However, ground-truth camera parameters are not always available in downstream tasks. To bridge this gap, [61, 18, 22] have proposed dedicated models to estimate camera intrinsics from monocular observations. In OpenVO, we employ the WildCamera [61] and Metric3Dv2 [16] to infer camera intrinsics and metric depth for 3D flow construction, guided by the 2D optical flow from MaskFlowNet [59].
Multi-Time-Scale Learning. Despite being one of the key quantities in temporally-driven problems, including video understanding, visual odometry, sequential decision making, and dynamical systems, observation rate (i.e., the time step size $\Delta t=1/f$) is typically fixed during the training process and not modeled explicitly by prior VO methods. Specifically, in frame-to-frame visual odometry, existing state-of-the-art models [21, 20, 47] take a pair of consecutive frames $(I_{t-1},I_{t})$, optionally along with auxiliary conditioning modality, such as language prior $c_{text}$, to estimate the ego motion. These models completely overlook the impacts of the temporal interval $\Delta t$ on the learning process, which might encode important temporal dynamics in visual odometry, such as velocity. As a result, they are trained under a single time scale (e.g., a fixed sampling rate of 20 Hz), leading to a subtle form of temporal overfitting [33] – where the learned motion representation is implicitly optimized for that specific sampling rate – and experiencing potential performance degradation when deployed under mismatched observation frequency (e.g., 12 Hz). Although these phenomena are underexplored in visual odometry and computer vision, prior studies in reinforcement learning [43] and world models [33] have shown that temporal overfitting to a fixed observation rate significantly degrades model robustness under unseen time scales. Motivated by these findings, we introduce a novel time-aware flow encoder in OpenVO and train the model across a diverse range of observation frequencies, exposing it to multiple temporal dynamic scales. This strategy effectively improves OpenVO’s accuracy across unseen large-scale autonomous driving benchmarks with varying frame rates.
@@ 3 Method
In the open world, dashcam videos are recorded across a vast range of devices with diverse lenses, focal lengths, and frame rates, posing significant challenges due to the absence of calibration information. When attempting to reconstruct camera trajectories from such Internet-scale footage, this lack of calibration becomes a fundamental obstacle. Traditional VO methods [34, 40, 6, 47, 42] assume access to known camera intrinsics to recover metric-scale motion, but this assumption breaks down in many downstream tasks, such as processing unconstrained online dashcam videos. We aim to develop a generalizable visual odometry system capable of extracting temporally consistent geometry from videos captured by uncalibrated cameras operating at arbitrary frame rates.
@@ 3.2 Geometry-Aware Context Encoder
which describes the direction of the 3D viewing ray passing through pixel $(u,v)$ in the camera coordinate system. This intrinsic ray field captures the capture the geometric projection pattern across the image plane: pixels near the optical center correspond to narrower ray cones, while those toward the periphery represent wider projection angles.
Depth tokenizer: We further incorporate metric depth to provide scene-scale information. Specifically, we leverage a pretrained metric depth estimator Metric3Dv2 [16] to obtain per-pixel depth $\mathbf{D}$. Each ray direction from Eq. 3 is then modulated by the corresponding depth value $\mathbf{M(u,v)}=\mathbf{D}(u,v)\cdot r(u,v)$. This operation projects each pixel onto its corresponding 3D point relative to the camera origin. The resulting spatial distribution $\mathbf{M(u,v)}$ effectively reconstructs the scene structure up to metric scale, capturing both the directionality and magnitude of each visual ray. Finally, we assemble the token set $\left[r,\mathbf{M},\mathbf{D}\right]$ representing a geometry-aware encoding of the camera projection behavior and scene structure. These tokens are processed by a Geometry-Aware Context Encoder composed of a stack of eight self-attention layers, which learn to correlate the intrinsic and depth priors with the surrounding visual context. This produces a unified geometric embedding that generalizes across uncalibrated dashcam sources and provides a consistent, metric-scale prior for world-coordinate egomotion estimation.
@@ 4.1 Experimental Setup
Evaluation metrics: We evaluate our method using three standard metrics: Translation Error ($t_{err}$), Rotation Error ($r_{err}$), and Absolute Trajectory Error (ATE), along with a recently introduced Scale Error ($s_{err}$) [20, 21]. The $t_{err}$ measures the average translational drift (in %) between the predicted and ground-truth trajectories, reflecting how accurately the model estimates displacement over time. The $r_{err}$ quantifies the angular deviation (in /100 m) between predicted and true orientations, indicating how well rotational motion is captured. The ATE assesses global trajectory consistency by aligning the predicted and ground-truth paths and computing the root-mean-square (RMS) positional difference, revealing how local errors accumulate over time. Finally, the $s_{err}$ measures the relative discrepancy in translation magnitudes, evaluating how faithfully the model preserves metric scale. It is computed as the mean difference between L2 norms of translation vectors across timesteps.
@@ 4.3 Ablation Studies
Analysis of varying temporal gap is presented in Tab. 4. To set up this configuration, for each dashcam video, we subsample the original framerate by a factor of $k$ that is, we skip (k1)(k-1) frames between consecutive inputs and apply the same subsampling to the corresponding ground-truth camera poses. The original capture rates are 10 Hz for KITTI, 12 Hz for nuScenes, and 20 Hz for Argoverse 2. We compare our OpenVO with the conventional ZeroVO, which disregards temporal frequency during training. As shown, ZeroVO exhibits substantial degradation when evaluated under varying frame rates (e.g., KITTI 2.5 Hz with 553.52 ATE; nuScenes 6 Hz with 21.55 ATE and Argoverse 2 20 Hz with 27.38 ATE), highlighting its inability to adapt to changes in temporal dynamics. In contrast, OpenVO consistently achieves lower translation, rotation, and trajectory errors across all frame rates by explicitly encoding frame-rate–aware temporal embeddings. This demonstrates that modeling temporal dynamics is critical for robust egomotion estimation under real-world video conditions, where frame rates are often inconsistent.
5 Discussion
In summary, we introduce OpenVO, a real-world–scale egomotion estimation framework that models temporal dynamics and geometric consistency from uncalibrated dashcam videos. Through extensive evaluations on KITTI, nuScenes, and Argoverse 2, OpenVO highlights the importance of time-aware, geometry-driven learning for scalable and robust visual odometry in real-world driving scenarios.
@@ 9 Qualitative Results
We provide additional qualitative demo videos in our supplementary materials!

### v2

@@ 2 Related Work
Geometric Prior Models. Geometric priors such as metric depth and camera intrinsics are crucial for 3D reconstruction and understanding [37, 36, 35] and improving scale consistency in visual odometry [23, 61, 47]. Monocular foundation metric depth models [62, 17, 4, 43, 42, 42, 59, 60, 3] enable absolute-scale depth prediction with strong zero-shot generalization across divese domains through large-scale training. Among these, some rely on known camera parameters [62, 17], while others remove this assumption and even achieve competitive performance in estimating focal lengths from monocular observation [4]. However, ground-truth camera parameters are not always available in downstream tasks. To bridge this gap, [68, 20, 24] have proposed dedicated models to estimate camera intrinsics from monocular observations. In OpenVO, we employ the WildCamera [68] and Metric3Dv2 [17] to infer camera intrinsics and metric depth for 3D flow construction, guided by the 2D optical flow from MaskFlowNet [66].
Multi-Time-Scale Adaptive Learning. Despite being one of the key quantities in temporally-driven problems, including video understanding, visual odometry, sequential decision making, and dynamical systems, observation rate (i.e., the time step size $\Delta t=1/f$) is typically fixed during the training process and not modeled explicitly by prior VO methods. Specifically, in frame-to-frame visual odometry, existing state-of-the-art models [23, 22, 53] take a pair of consecutive frames $(I_{t-1},I_{t})$, optionally along with auxiliary conditioning modality, such as language prior $c_{text}$, to estimate the ego motion. These models completely overlook the impacts of the temporal interval $\Delta t$ on the learning process, which might encode important temporal dynamics in visual odometry, such as velocity. As a result, they are trained under a single time scale (e.g., a fixed sampling rate of 20 Hz), leading to a subtle form of temporal overfitting [39] – where the learned motion representation is implicitly optimized for that specific sampling rate – and experiencing potential performance degradation when deployed under mismatched observation frequency (e.g., 12 Hz). Although these phenomena are underexplored in visual odometry, prior studies in object detection [38, 65], reinforcement learning [49] and world models [39] have shown that temporal overfitting to a fixed observation rate significantly degrades model robustness under unseen time scales. Motivated by these findings, we introduce a novel time-aware flow encoder and train the model across a diverse range of observation frequencies, exposing it to multiple temporal dynamic scales. This strategy effectively improves OpenVO’s accuracy across unseen large-scale autonomous driving benchmarks with varying frame rates.
@@ 3 Method
In the open world, dashcam videos are recorded across a vast range of devices with diverse lenses, focal lengths, and frame rates, posing significant challenges due to the absence of calibration information. When attempting to reconstruct camera trajectories from such Internet-scale footage, this lack of calibration becomes a fundamental obstacle. Traditional VO methods [40, 46, 6, 53, 48] assume access to known camera intrinsics to recover metric-scale motion, but this assumption breaks down in many downstream tasks, such as processing unconstrained online dashcam videos. We develop a generalizable VO system capable of extracting temporally metric-consistent geometry from videos captured by uncalibrated cameras with arbitrary frame rates.
@@ 3.2 Geometry-Aware Context Encoder
which describes the direction of the 3D viewing ray passing through pixel $(u,v)$ in the camera coordinate system. This intrinsic ray field captures the geometric projection pattern across the image plane: pixels near the optical center correspond to narrower ray cones, while those toward the periphery represent wider projection angles.
Depth tokenizer: We further incorporate metric depth to provide scene-scale information. Specifically, we leverage a pretrained metric depth estimator Metric3Dv2 [17] to obtain per-pixel depth $\mathbf{D}$. Each ray direction from Eq. 3 is then modulated by the corresponding depth value $\mathbf{M(u,v)}=\mathbf{D}(u,v)\cdot r(u,v)$. This operation projects each pixel onto its corresponding 3D point relative to the camera origin. The resulting spatial distribution $\mathbf{M(u,v)}$ effectively reconstructs the scene structure up to metric scale, capturing both the directionality and magnitude of each visual ray. Finally, we assemble the token set $\left[r,\mathbf{M},\mathbf{D}\right]$ representing a geometry-aware encoding of the camera projection behavior and scene structure. These tokens are processed by a Geometry-Aware Context Encoder composed of a stack of eight self-attention layers, which learn to correlate the intrinsic and depth priors with the surrounding visual context. This produces a unified geometric embedding that generalizes across uncalibrated dashcam sources, providing a consistent world-coordinate egomotion estimation.
@@ 4.1 Experimental Setup
Evaluation metrics: We evaluate our method using three standard metrics: Translation Error ($t_{err}$), Rotation Error ($r_{err}$), and Absolute Trajectory Error (ATE), along with a recently introduced Scale Error ($s_{err}$) [22, 23]. The $t_{err}$ measures the average translational drift (in %) between the predicted and ground-truth trajectories, reflecting how accurately the model estimates displacement over time. The $r_{err}$ quantifies the angular deviation (in /100 m) between predicted and true orientations, indicating how well rotational motion is captured. The ATE (in meters) assesses global trajectory consistency by aligning the predicted and ground-truth paths and computing the root-mean-square (RMS) positional difference, revealing how local errors accumulate over time. Finally, the $s_{err}$ measures the relative discrepancy in translation magnitudes, evaluating how faithfully the model preserves metric scale. It is computed as the mean difference between L2 norms of translation vectors across timesteps. Our metric & eval protocol follow ZeroVO, using only first-frame alignment, i.e., we set the first predicted pose to match the ground-truth pose and apply no SE(3) or Sim(3) alignment afterward. The models must predict metric-scale poses directly and no GT alignment applied for a fair comparison.
@@ 4.3 Ablation Studies
Analysis of varying temporal gap is presented in Tab. 4. To set up this configuration, for each dashcam video, we subsample the original framerate by a factor of $k$ that is, we skip (k1)(k-1) frames between consecutive inputs and apply the same subsampling to the corresponding ground-truth camera poses. The original capture rates are 10 Hz for KITTI, 12 Hz for nuScenes, and 20 Hz for Argoverse 2. We compare our OpenVO with the conventional ZeroVO. As shown, ZeroVO exhibits substantial degradation when evaluated under varying frame rates, highlighting its inability to adapt to changes in temporal dynamics. This demonstrates that modeling temporal dynamics is critical for robust egomotion estimation under real-world video conditions, where frame rates are often inconsistent.
5 Conclusion
We present OpenVO, a real-world–scale egomotion estimation framework that models temporal dynamics and geometric consistency from uncalibrated dashcam videos.
@@ 5 Conclusion
Acknowledgments: This project is supported in part by Barry Mersky and Capital One Endowed Professorships.
[…]
@@ 7 Analysis
Time-Conditioning Layers P.E. $t_{err}$ $r_{err}$ ATE
Cross-Attn (PE($\Delta t$) in Eq.1) 12.74 4.22 155.30
Cross-Attn (single scalar $\Delta t$) 14.74 4.78 180.16
Table 8: Time-Conditioning Layers design. Using Cross-Attention layers as alternatives to the design of Time-Conditioning Layers and the use of our proposed positional encoding (P.E.).
Ablation P.E. Flow Geometry $t_{err}$ $r_{err}$ ATE
OpenVO 17.12 4.42 158.67
OpenVO - - -
OpenVO 9.59 3.54 117.50
Table 9: Time-Conditioning Layers design. Using Cross-Attention layers as alternatives to the design of Time-Conditioning Layers and the use of our proposed positional encoding (P.E.).
Methods $t_{err}$ $r_{err}$ ATE
PnP+Metric3Dv2+MaskFlow 54.38 2.81 302.74
DroidCalib [15]: no Sim(3) (fair) 69.16 0.625 382.65
DroidCalib [15]: w/ Sim(3) (unfair) 11.01 0.625 73.18
ViPE [18]: no Sim(3) (fair) 13.29 0.721 137.15
ViPE [18]: w/ Sim(3) (unfair) 10.15 0.721 72.73
Table 10: Geometry-based and optimization-based methods.
Motivation: $\Delta t$ is a conditioning signal indicating the temporal stride of the input pair, serving as a proxy for providing temporal context so the network can produce stable pose increments across different sampling intervals as shown in Tab. 4 (main paper). The purpose of conditioning is not to deterministically scale the predicted pose or to “force” a particular motion magnitude, but to adapt to a wide range of motion magnitudes – a behavior that prior methods cannot achieve. Without TCL, the model must self-discover motion patterns without any explicit temporal modeling, leading to suboptimal learning as the single flow encoder backbone is overloaded with multiple conflicting objectives (differentiating small, medium, and large motion magnitudes). Our TCL with the proposed P.E. provides the necessary temporal bridge, stabilizing and improving learning. Unlike prior work, our Context Encoder jointly encodes image–depth inputs and camera parameters (L.361), enabling the model to be camera-aware and generalize effectively to any unseen camera settings during testing.
Ablation on Time-Conditioning Layers: We further replace Cross-Attention as the Time-Conditioning Layer (TCL) and ablate our positional encoding in Tab. 8. When disregard our proposed P.E., we use a single scalar $\Delta t$ as a temporal condition. The results validate the effectiveness of our design, potentially motivating the 3D/4D reconstruction communities to leverage temporal dynamics for computer vision problems.
Ablation on the proposed components: We provide an ablation on the feature stream of OpenVO in Tab. 9. Without the flow branch, we cannot estimate the geometry difference between two consecutive frames. OpenVO without explicit geometry (depth) modeling and camera information suffers from generalization with $t_{err}$ of 17.12.
Comparison with geometric-based and optimization-based methods: Traditional closed-world monocular VO/SLAM approaches with known intrinsics, backend optimization & loop closure, are evaluated with known GT poses alignment. Our OpenVO is evaluated under a strictly harder setting: unknown intrinsics, no GT alignment, no loop closure, and zero-shot cross-dataset generalization. We showcase some state-of-the-art (2026-March) and traditional geometry-based methods on visual odometry in Tab. 10. Our OpenVO still achieves state-of-the-art results on reconstructing real-world scale egomotion without the need of alignment.